OpenStudy (avai70178):

How do I solve this absolute value equation? -10-|10n+8|=-78

OpenStudy (anonymous):

-10-(10n+8)=-78 and by switching the signs of everything -[-10-(10n+8)]=-78

OpenStudy (avai70178):

okay but what happens with the negative in front of the absolute value?

OpenStudy (anonymous):

I mean switching the signs of the side with variable in it.

zepdrix (zepdrix):

\[\large\rm -10-|10n+8|=-78\]Should probably add 10 to each side before dealing with the signs that the absolute value gives us :)\[\large\rm -|10n+8|=-68\]

OpenStudy (anonymous):

You make it into parentheses, so just ignore the absolute value Negative times negative makes it positive

OpenStudy (avai70178):

oh okay i see. Thank you so much! :)

OpenStudy (anonymous):

Did you get your first and second solution?

OpenStudy (avai70178):

not yet one sec

OpenStudy (avai70178):

10n+8=+or- 68 right so far?

OpenStudy (anonymous):

Yea that is correct.

OpenStudy (avai70178):

i got n=-38/5 and n=6

OpenStudy (anonymous):

the -68 involves a fraction.

OpenStudy (anonymous):

Yea, I got that too.

OpenStudy (anonymous):

Nice job!

OpenStudy (avai70178):

Thanks! You really helped! :D

OpenStudy (anonymous):

No problem!